# Asymmetric Contrastive Multimodal Learning for Advancing Chemical Understanding

### v3

@@ Abstract
The versatility of multimodal deep learning holds tremendous promise for advancing scientific research and practical applications. As this field continues to evolve, the collective power of cross-modal analysis promises to drive transformative innovations, leading us to new frontiers in chemical understanding and discovery. Hence, we introduce Asymmetric Contrastive Multimodal Learning (ACML) as a novel approach tailored for molecules, showcasing its potential to advance the field of chemistry. ACML harnesses the power of effective asymmetric contrastive learning to seamlessly transfer information from various chemical modalities to molecular graph representations. By combining pre-trained chemical unimodal encoders and a shallow-designed graph encoder, ACML facilitates the assimilation of coordinated chemical semantics from different modalities, leading to comprehensive representation learning with efficient training. We demonstrate the effectiveness of this framework through large-scale cross-modality retrieval and isomer discrimination tasks. Additionally, ACML enhances interpretability by revealing chemical semantics in graph presentations and bolsters the expressive power of graph neural networks, as evidenced by improved performance in molecular property prediction tasks from MoleculeNet. ACML exhibits its capability to revolutionize chemical research and applications, providing a deeper understanding of the chemical semantics of different modalities.
@@ 2.1 ACML framework
The proposed ACML framework is built upon the contrastive learning framework inspired by [13]. It effectively promotes correspondences between modalities of the same molecule (positive pairs) while contrasting them with different molecules (negative pairs). The whole pipeline, elaborated in Section 4.1, is composed of four components: an unimodal encoder for molecular modalities, a graph encoder formalized by graph neural networks, a non-linear projection head, and the temperature-scaled cross-entropy contrastive loss.
@@ 2.2 Cross-modality retrieval
Molecules can be represented by a variety of modalities, such as depiction images, SMILES notations, and traditional chemical characterization techniques. However, each modality possesses distinct capabilities and constraints when it comes to deciphering molecular structures. In the context of structural attributes, molecular depiction image and isomeric SMILES notation offer a detailed overview of the whole molecule from atom level to molecule level, facilitating the elucidation of their three-dimensional atomic configurations. NMR spectra provide distinct windows into the carbon and hydrogen atomic landscapes of the molecule, offering partial structural information about the molecule. Mass spectra provide crucial information about molecular mass and patterns of fragmentation, encompassing both molecular-level and motif-level information. To evaluate each ACML model, we conduct a cross-modality retrieval task: determine whether the model can accurately match a chemical modality to its corresponding graph modality from a large database of molecular graphs. Higher retrieval accuracy indicates better multi-modal coordination between graph and chemical representations. Note that the sample in the graph database never appears in the training phase.
@@
Figure 2: Cross-modality retrieval accuracy for all ACML models, including G-SMILES, G-image, G-1H NMR, G-13C NMR, G-GCMS, and G-LCMS. In each ACML model. The X-axis represents the molecular pool ranging from 1000 (1K) to 1 million (1M) molecules. From this pool, the model is tasked to retrieve the target molecule by ranking the similarity scores. The Y-axis is the top k accuracy, which measures whether the target molecule is within the most similar k molecules the model recommended. All of our models can successfully identify the target molecule from more than 5000 candidates. Even when the molecular pool is as large as 1 million, models trained using the image and SMILES representations still achieve more than 90% accuracy.
@@
Figure 2 presents the retrieval accuracy results for all ACML models, including G-SMILES, G-image, G-1H NMR, G-13C NMR, G-GCMS, and G-LCMS. These results are obtained by calculating the top k accuracy when matching one molecule’s chemical modality to its corresponding graph within a graph database of varying scales. The task becomes more challenging as the scale increases. When the scale transitions from 1 thousand (K) to 1 million (M), the G-Image gives the best results: the top 10 accuracy drops from 100% to 95.0%, which only experiences a reduction of 5.0%. G-SMILES gives the second best results, as top 10 accuracy drops from 100% to 87.5%. When it comes to G-1H NMR, the results are not as good as G-Image and G-SMILES, as expected. Specifically, when the scale transitions from 1K to 10K, the top 10 accuracy drops from 96.1% to 84.2%. And with 1M unseen molecules, the top 10 accuracy is 38.7%. G-13C NMR spectrum exhibits similar performance, as the top 10 accuracy from 96.3% to 83.9% until 43.4% when the scale transitions from 1K to 10K and up to 1M. Notably, G-13C NMR performs better than G-1H NMR when the search space expands. Intuitively, this aligns with the expectation, since the 13C NMR peaks are typically sharp and barely overlap in small molecules. On the contrary, the 1H NMR spectrums are more prone to signal overlap, especially in complex molecular structures. Although the encoders of 13C NMR and 1H NMR have been designed to cope with such differences, the training dataset inevitably falls short of encompassing the entirety of structural variations within the zero-shot validation set. Additionally, 1H NMR might encounter sensitivity issues, especially for compounds with low hydrogen content. Nonetheless, both the G-1H NMR and G-13C NMR exhibit satisfactory discrimination power in identifying the most probable graphs given the NMR sequences. This demonstrates the flexibility of our ACML framework with features from different spaces and dimensions generated from different designs of encoders.
@@
When it comes to mass spectra modality, the results do not align as favorably as with images, smiles, and NMR spectra, as initially anticipated. For G-GCMS, when the scale transitions from 1K to 10k, the top 10 accuracy experiences a certain degree of reduction from 70.5% to 39.5%. When the validation database extends to 1M, the top 10 accuracy dropped to 7.8%. For G-LCMS, when the scale transitions from 1K to 10K, the top 10 accuracy experiences a certain degree of reduction from 32.4% to 14.9%. When the validation database extends to 1M, the top 10 accuracy dropped to 1.7%. One may observe that G-GCMS has better performance on molecular recognition than G-LCMS. The reason for that is GCMS is typically carried out under harsh conditions, resulting in a wide array of fragmentation that unveils the molecular composition. In contrast, LCMS is generally performed under milder conditions, resulting in reduced fragmentation. Particularly, LC-MS/MS primarily concentrates on localized information near the precise molecular weight. However, in terms of cross-modality retrieval, LCMS offers less comprehensive information.
@@ 2.3 Isomer discrimination
Besides calculating accuracy among a large group of molecules, in which many molecules could be intrinsically different, isomer recognition stands as one of the most challenging tasks in molecular identification due to the intricacy of isomeric structures. Distinguishing between stereoisomers, which share the same connectivity but differ in spatial arrangement, poses a particular challenge and demands advanced analytical methods and meticulous data interpretation. Case studies have been conducted to demonstrate the capabilities of NMR modalities in terms of isomer recognition. From the validation set, several pairs of complex isomers have been picked up randomly. These isomers not only have identical formulas but also feature highly similar functional groups. In this investigation, both 13C NMR and 1H NMR yield comparable results and effectively distinguish among the majority of instances. For each scenario, the ACML model is supplied with sets of NMRs, alongside confidence scores indicating the likelihood of these NMRs aligning with the corresponding pair of graphs. A greater confidence score signifies a stronger alignment according to our proposed framework. Four cases of 13C NMR are shown in Figure 3, where for each case, interested pairs of 13C NMR spectrum and possible molecular graphs are presented to the model. The model selects the molecule of the highest possibility to be the corresponding decipher of the 13C NMR spectrum. The proposed model correctly identifies 3 isomers out of 4. Similarly, four cases of 1H NMR are shown in Figure 4, where for each case, interested pairs of 1H NMR spectrum and possible molecular graphs are presented to the model. The proposed model correctly identifies 3 isomers out of 4.
@@ 2.4 Visualization of molecular representations
In the ACML framework, the graph modality serves as an information receptor capable of bringing forth the information or chemical rules from another modality via coordination. Thus, it may reveal distinct dimensions of chemical significance among diverse chemical modalities. To illuminate variations of chemical semantics embedded in hidden graph representations acquired from various ACML frameworks, we computed graph embeddings of previously unseen 30,000 molecules using pre-trained graph encoders for analysis. Given 8 important chemical characteristics in drug discovery: molecular weight (MW), the logarithmic partition coefficient (LogP), hydrogen-bonding acceptor (HBA), hydrogen-bonding donor (HBD), topological polar surface area (TPSA), the number of rotatable bonds (#R-Bonds), and the number of stereogenic (chiral) centers (#C-Centers), we decided to examine how these characteristics correlate with the graph embeddings. The scope of this inquiry extends to understanding the impact of these modalities on crucial parameters within the domain of drug discovery, highlighting the strengths of each molecular modality.
@@
We decomposed the feature dimensions via PCA [48] and mapped the resulting embeddings onto a two-dimensional plane. The outcomes were visualized as 2D points with coloring based on their corresponding chemical property values. Figure 5 presents the visualization results. Notably, a remarkable consistency in the distribution of points concerning the chemical property values was observed. This observation underscores the proficiency of the pre-trained graph encoder in capturing the inherent chemical semantics. We then conducted a quantitative evaluation of the relationship between graph embedding and the chemical properties using Pearson correlation coefficients (PCC) [49]. In our approach, we first established a linear regression model to determine the optimal linear combination of two PCA components that maximizes the Pearson correlation score between the linear combination of PCA components and chemical properties of interest. Table 1 presents the PCC scores of all ACML frameworks.
@@
The knowledge of the chemical modalities was captured and highlighted by the degree of correlation, particularly with respect to key properties relevant to drug discovery. Regarding MW, G-Image achieved the highest score, followed closely by G-GCMS. Molecular weights can effectively be expressed by the pixels of molecular depiction image, and extracted from mass spectrometry data provided by GCMS. This accounts for the superior performance of these two models in this context. For HBA, G-SMILES emerged as the top performer. It is expected because electronegative HBA atoms, such as oxygen (O) or nitrogen (N), are directly denoted in SMILES. For HBD, G-Image demonstrates the best performances, as HBD groups, such as hydroxyl groups (-OH), amine groups (-NH2), and carboxylic acids (-COOH), are explicitly illustrated in the images. However, hydrogens are not always contained in SMILES, resulting in G-SMILES has a slightly lower performance. In the case of LogP, G-13C NMR and G-1H NMR are the game winners. This outcome is attributable to the fact that LogP measures the solubility of a solute in both water and organic solvents, making it highly reliant on electronic information within the molecule. This electronic information is well captured by the two NMR techniques, and it remains concealed in other modalities such as Image, SMILES, or mass spectrometry. For #R-Bonds, G-1H NMR exhibited the strongest performance because the rotation of chemical bonds can be derived from the magnetic equivalence of vicinal and geminal protons. When considering #C-Centers, G-SMILES achieves the highest correlation coefficient of 0.8297. This strong correlation can be elucidated by the inherent advantage of SMILES: atoms with chiral centers are directly annotated in SMILES representations, that is, “@" represents chirality with inward direction and “@@" represents chirality with outward direction.
@@
Remark 2. These 8 chemical properties are very important in drug discovery. If the graph neural network is able to reveal this information, it offers the potential for deep models to understand the basic philosophy of molecular design and would even make breakthroughs in drug discovery. Note that these properties are never seen in ACML learning and the molecules we selected are also never seen in training, the surprising findings demonstrate our ACML framework can make a transition of chemical meaning to graph encoder, enhancing its interpretability.
Figure 5: Graph embedding visualization via PCA.
@@
G-Image 0.742 0.672 0.724 0.355 0.736 0.325 0.630 0.597

### v4

@@ Abstract
The versatility of multimodal deep learning holds tremendous promise for advancing scientific research and practical applications. As this field continues to evolve, the collective power of cross-modal analysis promises to drive transformative innovations, leading us to new frontiers in chemical understanding and discovery. Hence, we introduce Asymmetric Contrastive Multimodal Learning (ACML) as a novel approach tailored for molecules, showcasing its potential to advance the field of chemistry. ACML harnesses the power of effective asymmetric contrastive learning to seamlessly transfer information from various chemical modalities to molecular graph representations. By combining pre-trained chemical unimodal encoders and a shallow-designed graph encoder with 5 layers, ACML facilitates the assimilation of coordinated chemical semantics from different modalities, leading to comprehensive representation learning with efficient training. We demonstrate the effectiveness of this framework through large-scale cross-modality retrieval and isomer discrimination tasks. Additionally, ACML enhances interpretability by revealing chemical semantics in graph presentations and bolsters the expressive power of graph neural networks, as evidenced by improved performance in molecular property prediction tasks from MoleculeNet and TDC. ACML exhibits its capability to revolutionize chemical research and applications, providing a deeper understanding of the chemical semantics of different modalities.
@@ 2.1 ACML Framework
The proposed ACML framework is built upon the contrastive learning framework inspired by [13]. It effectively promotes correspondences between modalities of the same molecule (positive pairs) while contrasting them with different molecules (negative pairs). The whole pipeline, elaborated in Section A.1, is composed of four components: an unimodal encoder for molecular modalities, a graph encoder formalized by graph neural networks, a non-linear projection head, and the temperature-scaled cross-entropy contrastive loss.
@@ 2.2 Cross-modality Retrieval
Molecules can be represented by a variety of modalities, such as depiction images, SMILES notations, and traditional chemical characterization techniques. However, each modality possesses distinct capabilities and constraints when it comes to deciphering molecular structures. In the context of structural attributes, molecular depiction image and isomeric SMILES notation offer a detailed overview of the whole molecule from atom level to molecule level, facilitating the elucidation of their three-dimensional atomic configurations. NMR spectra provide distinct windows into the carbon and hydrogen atomic landscapes of the molecule, offering partial structural information about the molecule. Mass spectra provide crucial information about molecular mass and patterns of fragmentation, encompassing both molecular-level and motif-level information. To evaluate each ACML model, we conduct a cross-modality retrieval task: determine whether the model can accurately match a chemical modality to its corresponding graph modality from a large database of molecular graphs. Higher retrieval accuracy indicates better multi-modal coordination between graph and chemical representations. Note that the sample in the graph database never appears in training.
@@
Figure 2: Cross-modality retrieval accuracy across all ACML models, including G-SMILES, G-Image, G-1H NMR, G-13C NMR, G-GCMS, and G-LCMS. The x-axis represents the molecular pool ranging from 1,000 (1K) to 1 million (1M) molecules on the log scale. Top-10 and top-100 accuracy of random guesses are incorporated. Each model retrieves the target molecule by ranking similarity scores within this pool. The y-axis is the top k accuracy, measuring whether the target molecule is within the k most similar molecules recommended by the model. All models consistently identify the target molecule from pools exceeding 5,000 candidates. Notably, Image- and SMILES-based models maintain over 90% top-100 accuracy even with a pool of 1M molecules.
@@
Figure 2 presents the retrieval accuracy results for all ACML models, including G-SMILES, G-Image, G-1H NMR, G-13C NMR, G-GCMS, and G-LCMS. These results are obtained by calculating the top k accuracy when matching one molecule’s chemical modality to its corresponding graph within a graph database of varying scales. The task becomes more challenging as the scale increases. When the scale transitions from 1,000 (1K) to 1 million (1M), the G-Image gives the best results: the top-10 accuracy drops from 100% to 94.8%, which only experiences a reduction of 5.2%. G-SMILES gives the second best results, as top 10 accuracy drops from 100% to 86.5%. When it comes to G-1H NMR, the results are not as good as G-Image and G-SMILES, as expected. Specifically, when the scale transitions from 1K to 10K, the top-10 accuracy drops from 94.9% to 79.7%. And with 1M unseen molecules, the top-10 accuracy is 43.6%. G-13C NMR spectrum exhibits similar performance, as the top-10 accuracy from 93.9% to 81.2% until 43.8% when the scale transitions from 1K to 10K and up to 1M. Notably, G-13C NMR performs better than G-1H NMR when the search space expands. Intuitively, this aligns with the expectation, since the 13C NMR peaks are typically sharp and barely overlap in small molecules. On the contrary, the 1H NMR spectrums are more prone to signal overlap, especially in complex molecular structures. Although the encoders of 13C NMR and 1H NMR have been designed to cope with such differences, the training dataset inevitably falls short of encompassing the entirety of structural variations within the zero-shot validation set. Additionally, 1H NMR might encounter sensitivity issues, especially for compounds with low hydrogen content. Nonetheless, both the G-1H NMR and G-13C NMR exhibit satisfactory discrimination power in identifying the most probable graphs given the NMR sequences. This demonstrates the flexibility of our ACML framework with features from different spaces and dimensions generated from different designs of encoders.
@@
When it comes to mass spectra modality, the results do not align as favorably as with Images, SMILES, and NMR spectra, as initially anticipated. For G-GCMS, when the scale transitions from 1K to 10K, the top-10 accuracy experiences a certain degree of reduction from 75.2% to 46.3%. When the validation database extends to 1M, the top-10 accuracy dropped to 19.4%. For G-LCMS, when the scale transitions from 1K to 10K, the top 10 accuracy experiences a certain degree of reduction from 45.5% to 32.2%. When the validation database extends to 1M, the top-10 accuracy dropped to 16.9%. One may observe that G-GCMS has better performance on molecular recognition than G-LCMS. The reason for that is GCMS is typically carried out under harsh conditions, resulting in a wide array of fragmentation that unveils the molecular composition. In contrast, LCMS is generally performed under milder conditions, resulting in reduced fragmentation. Particularly, LC-MS/MS primarily concentrates on localized information near the precise molecular weight. However, in terms of cross-modality retrieval, LCMS offers less comprehensive information.
@@ 2.3 Isomer Discrimination
Task Description.
Besides calculating accuracy among a large group of molecules, in which many molecules could be intrinsically different, isomer recognition stands as one of the most challenging tasks in molecular identification due to the intricacy of isomeric structures. Distinguishing between stereoisomers, which share the same connectivity but differ in spatial arrangement, poses a particular challenge and demands advanced analytical methods and meticulous data interpretation.
To thoroughly evaluate the model’s discriminative power, we selected 140 isomer pairs from the test dataset in a systematic way, including both structural isomers and stereoisomers. First, all SMILES representations were converted to molecular formulas using the RDKit package in Python. Molecules sharing the same molecular formula were grouped, and pairs within each group were identified as isomers. To ensure meaningful comparisons, we visualized the isomer pairs and excluded those with significantly different structures (e.g., ring vs. chain) while retaining only pairs with high structural similarity (differing by fewer than three bonds). From the remaining isomer pairs, we randomly selected 40 stereoisomers and 100 structural isomers to form a representative test set. The selected pairs are provided in Supplementary Material S1, with the isomer distribution shown in Table 1.
Results and Analysis.
Both human experts and the ACML model were tasked with associating NMR spectra to the most likely molecular isomers. Besides the binary classifications, the ACML model produces confidence scores that indicate the alignment strength between the NMR spectra and potential molecular structures. Higher confidence scores represent stronger alignment using our proposed framework. When comparing the expert’s performance across different NMR spectroscopy techniques, the expert demonstrated better accuracy using HNMR compared to CNMR spectra. This is consistent with the fact that protons have the highest natural abundance ($\sim$ 99%), resulting in denser HNMR signals that provide more information for distinguishing between isomers. In contrast, 13C’s lower natural abundance ($\sim$ 1%) means fewer signals are available for analysis. The performance comparison between the model and the expert is presented in Table 2. Our model outperforms the human expert for both 1H NMR and 13C NMR tasks, with a much more significant lead for 13C NMR tasks. This demonstrates the effectiveness of representation learning using our proposed framework.
Table 1: Distribution and Similarity of isomer pairs. Tanimoto Similarity [48] is used to measure the similarity between chemical structures.
Structural Isomer Stereoisomer
Count (Avg. Similarity) Count (Avg. Similarity)
1H NMR 47 (0.43) 15 (1.00)
13C NMR 61 (0.47) 17 (1.00)
Table 2: Performance Comparison between the Model and the Expert.
1H NMR 13C NMR
Structural Isomer Stereoisomer Structural Isomer Stereoisomer
Model Accuracy 82.6% 51.6% 82.9% 60.0%
Human Accuracy 78.0% 51.6% 51.9% 57.1%
To visualize the model’s performance, four cases of 13C NMR are shown in Figure 3 and 1H NMR examples are shown in Figure 4. The percentage on the arrows represents the ACML model’s confidence score, with higher percentages indicating greater certainty in matching the NMR spectra to specific molecular isomers. These scores provide a probabilistic measure of the model’s confidence in structural assignments.
@@ 2.4 Visualization of Molecular Representations
In the ACML framework, the graph modality serves as an information receptor capable of bringing forth the information or chemical rules from another modality via coordination. Thus, it may reveal distinct dimensions of chemical significance among diverse chemical modalities. To illuminate variations of chemical semantics embedded in hidden graph representations acquired from various ACML frameworks, we computed graph embeddings of previously unseen 30,000 molecules using pre-trained graph encoders for analysis. Given 8 important chemical characteristics in drug discovery: molecular weight (MW), the logarithmic partition coefficient (LogP), hydrogen-bonding acceptor (HBA), hydrogen-bonding donor (HBD), topological polar surface area (PSA), the number of rotatable bonds (#R-Bonds), and the number of stereogenic (chiral) centers (#C-Centers), we decided to examine how these characteristics correlate with the graph embeddings. The scope of this inquiry extends to understanding the impact of these modalities on crucial parameters within the domain of drug discovery, highlighting the strengths of each molecular modality.
Evaluations and Baselines.
@@
We decomposed the feature dimensions via PCA [49] and mapped the resulting embeddings onto a two-dimensional plane. The outcomes were visualized as 2D points with coloring based on their corresponding chemical property values. Figure 5 presents the visualization results. Notably, a remarkable consistency in the distribution of points concerning the chemical property values was observed. This observation underscores the proficiency of the pre-trained graph encoder in capturing the inherent chemical semantics. We then conducted a quantitative evaluation of the relationship between graph embedding and the chemical properties using Pearson correlation coefficients (PCC) [50]. In our approach, we first established a linear regression model to determine the optimal linear combination of two PCA components that maximizes the Pearson correlation score between the linear combination of PCA components and chemical properties of interest. We also showed the graph embeddings from other pretraining graph encoders and performed PCA in the same fashion to compare how the latent representations differ with respect to ACML. Among the baselines, L2P-GNN [51] and MGSSL [52] develop motif-level self-supervised tasks during pretraining, while MolCLR [30] and MoMu [53] use contrastive learning in pretraining.
Results and Analysis.
@@
Table 3 presents the PCC scores of all ACML frameworks and baselines. All graph embedding visualizations are provided in Supplementary Material S2. The knowledge of the chemical modalities was captured and highlighted by the degree of correlation, particularly with respect to key properties relevant to drug discovery. Among the methods analyzed, the ACML frameworks generally outperform baseline models, achieving the highest PCCs in seven out of eight properties, with the exception of #R-Bonds, where MoMu achieves the top PCC (0.617). Regarding MW, G-Image and G-SMILES are leading performers. Molecular weights can effectively be expressed by the pixels of molecular depiction image, and structural motifs encoded by SMILES. These account for the superior performance of these two models in this context. Additionally, G-GCMS and G-LCMS also demonstrate competitive performances. This can be attributed to their direct measurement of molecular mass in spectrometry. Regarding #HBA, G-SMILES emerges as the top performer. It is expected because electronegative HBA atoms, such as oxygen (O) or nitrogen (N), are directly denoted in SMILES. Regarding #HBD, G-Image is the top-performing model, as HBD groups, such as hydroxyl groups (-OH), amine groups (-NH2), and carboxylic acids (-COOH), are explicitly illustrated in the images. However, hydrogens are not always contained in SMILES, resulting in G-SMILES has a slightly lower performance. Regarding LogP, G-13C NMR is the game winners, followed by the MGSSL (BFS) baseline. This outcome is attributable to the fact that LogP measures the solubility of a solute in both water and organic solvents, making it highly reliant on electronic information within the molecule. This electronic information is well captured by the 13C NMR techniques, and it remains concealed in other modalities such as Image, SMILES, or mass spectrometry. Regarding PSA, G-SMILES achieves the highest, followed closely by G-Image. The strong performance of G-SMILES reflects its explicit representation of electronegative atoms, such as oxygen and nitrogen, which directly contribute to PSA. Besides, G-Image captures polar regions through pixel-based molecular depictions. Regarding #R-Bonds, MoMu demonstrates the highest PCC, while G-1H NMR closely follows as the second best and significantly outperforms all other modalities. The reason is the rotation of chemical bonds can be derived from the magnetic equivalence of vicinal and geminal protons, while not revealed in other modalities.
@@
Remark 2. These 8 chemical properties play vital roles in drug discovery. If the graph neural network is able to reveal this information, it offers the potential for deep models to understand the basic philosophy of molecular design and would even make breakthroughs in drug discovery. Note that these properties are never seen in ACML learning and the molecules we selected are also never seen in training, the surprising findings demonstrate our ACML framework can make a transition of chemical meaning to graph encoder, enhancing its interpretability.
Figure 5: Graph embedding visualization via PCA. For each property, one representative figure is shown using the ACML instantiation that achieves the highest PCC score.
@@
L2P-GNN 0.053 0.026 0.027 0.029 0.026 0.024 0.237 0.056
MGSSL (BFS) 0.259 0.639 0.618 0.700 0.670 0.331 0.789 0.371
MGSSL (DFS) 0.255 0.638 0.504 0.673 0.649 0.266 0.759 0.381
MoMu 0.263 0.305 0.225 0.386 0.280 0.617 0.609 0.213
MolCLR 0.228 0.294 0.207 0.164 0.244 0.378 0.558 0.182
G-Image 0.742 0.671 0.724 0.355 0.736 0.325 0.630 0.597